OpenStudy (anonymous):

m/-6 +6=41 need a soultion

jimthompson5910 (jim_thompson5910):

m/-6 +6=41 m +6(-6)=41(-6) .. Note: multiply EVERY term by the LCD -6 to clear out the fraction m-36 = -246 m = -246+36 m = -210